How do I solve this equation, "y/16+84=92"
Answer by lenny460(1073)

How do I solve this equation, "y/16+84=92"

Multiply each item by 16.
y/16 * 16 + 84 * 16 = 92 * 16
y/16 * 16 = y
84 * 16 = 1344
92 * 16 = 1472
We now have:
y + 1344 = 1472
Subtract 1344 from each side of the equation.
y + 1344 - 1344 = 1472 - 1344
Combine like terms:
1344 - 1344 = 0
1472 - 1344 = 128
y = 128
Answer:
y = 128








Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com